OpenStudy (anonymous):

f(x)=x^4-25x+144, x=1 Find the slope of the tangent line to graph the given value of x

OpenStudy (anonymous):

you need to find f'(1), or the derivative of f(x) evaluated at 1. f'(x) = 4x^3 - 25 f'(1) = 4-25 = -21 the slope of the tangent line (which is the slope of the line) at x=1 is -21.

OpenStudy (anonymous):

im soo sorry thats thhe wrong problem its : f(x)=x^4-25x^2+144, x=1 Find the slope of the tangent line to graph the given value of x

OpenStudy (anonymous):

f'(x) = 4x^3 - 50x f'(1) = 4 - 50 =-46

OpenStudy (anonymous):

is -46 y or is it the slope?

OpenStudy (anonymous):

its the slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but where y?

OpenStudy (anonymous):

umm, so i gave you the slope to the tangent line, would you like the entire equation of the tangent line? at x = 1, y is... f(1) = 4 - 25 + 144 = 123 the tangent line is of the form y = m*x + b, where m equals the slope, -46. now plug in y = 123 at x = 1 to get b... 123 = -46*1 + b b = 169 the equation of the tangent line is y = -46*x + 169 and y at x = 1 is 123 if you wanted that too

OpenStudy (anonymous):

tht should be 120 shouldnt it be?

OpenStudy (anonymous):

(1)^4-25(1)^2+144=120

OpenStudy (anonymous):

oh yeah, sorry

OpenStudy (anonymous):

its cool